Henry invested 12,000 in a new restaurant. When the restaurant was sold two years later, he recieved 27,000. Find his average annual return by solving the equation 12,000(1+r)^2= 27,000.
Answer by user_dude2008(1862)

12,000(1+r)^2= 27,000

(1+r)^2= 27000/12000


(1+r)^2=9/4

1+r=sqrt(9/4)

1+r=3/2

r=3/2-1

r=1/2

r=0.5